OpenStudy (anonymous):

Which of the following is a solution for the system y = -3x + 5 and y = 2x + 5? A. (1, 7) B. (5, 0) C. (2, -1) D. (0, 5)

OpenStudy (anonymous):

lets sub y = -3x + 5 into the second equation y = 2x + 5 -3x + 5 = 2x + 5 bring -3x over to the right and the positive five to the left. 0 = 5x x = 0 sub x = 0 into the first or second equation, lets use the second one. y = 2(0) + 5 y = 0 + 5 y = 5 so now we have x = 0 and y = 5 so the solution is (0,5)